if the fixed price of land is 1000 and firm pay 750 to each worker then calculate, Fixed cost, variable Cost, Marginal Cost, Average variable Cost (AVC), Average Fixed Cost(AFC), and Average total cost(ATC).



1
Expert's answer
2021-05-11T15:03:44-0400

labor           total product

0                0

 1               8

2               18

3               25

4              30

5               33

6               34



fixed cost is1000.

variable cost(VC)"=Wage\space of\space labor\space ×\space Quantity"

The marginal cost  (MC)"=\frac{Change\space in \space TC}

{Change \space in\space quantity}"

When the quantity changes from 8 to 18, the MC would be:

"MC=\frac{2500-1750}{18-8}=\frac{750}{10}=75"

Similarly, the MC at all levels can be calculated.

Total costs (TC) are the sum of fixed and variable costs.

The variable cost divided by quantity gives average variable cost (AVC). 

The formula for AVC is:"AVC=\frac{VC}{Q}"

For example: the AVC when the quantity is 18 is calculated:

"AVC=\frac{1500}{18}=83.33" The fixed cost divided by quantity gives average fixed cost (AFC).

The formula for AFC is: "AFC=\frac{FC}

Q"

The total cost divided by the quantity produced gives average total cost (ATC).

The formula for ATC is: "ATC=\frac{TC}{Q}"

The sum of AVC and AFC will also give ATC.

 The values so calculated are given in table below: